You have two Points, A and B, with the respective number pairs (10, 15) and (7, 21). What is the value of the slope of a line drawn between these two points?
Answer #1

Given two points - A (10,15) and B ( 7, 21 )

In order to find the slope of a line joining the given points, we need to folllw the following steps-

STEP-1

We first need to denote the given points in the following notations-

A ( x1, y1 ) = A ( 10, 15 ) , where x1 = 10 and y1= 15

B ( x2 , y2 ) = B ( 7,21) , where x2 = 7 and y2 = 21

STEP-2

Our next step involves calculating the slope of the line joining the two points.

We can calculate the slope by using the following formula-

Slope = ( y2 - y1 ) / ( x 2 - x1 )

or Slope = (21-15)/(7-10) = (6)/(-3) = -2

Thus, SLOPE = -2.

The slope indicates the direction of the line joining the two points. Since the slope is negative, it implies that the line is downward sloping.